I'm having trouble determining how to find delta T and the freezing point depression for Solutions #1 and #2. This is all the information I have. Please show all work. Thank you.

Mass of test tube 24.7795 g
Mass of tube + BZP 34.2490 g
Mass of BZP used 9.4695 g
Freezing point of pure BZP -9.80 kg/mL
Solution 1 Solution 2
Mass of container and Sample 11.7366 g 11.7366 g
Mass of container and Sample (after weighing) 11.0103 g 9.6156 g
Mass of unknown used 0.7263 g 2.1210 g
Freezing Point of solution 41.2 (C) 39.8 (C)
Freezing Point depression (delta T)
Molality of unknown
Molar Mass of unknown

Average molar mass of unknown

Freezing point of BZP is listed above...-9.80 C kg/Mol


the freezing point of pure BZP was 45.6 celsius
Answer #1

freezing point of pure BZP was 45.6 celsius

freezing point of solution was 41.2 celsius

hence depresssion in freezing point deltaT = |41.2 - 45.6| = 4.4 celsius (solution 1)

for solution 2 , 39.8 - 45.6 = 5.8 celsius

deltaT = Kf * m

where for water Kf = 1.853

hence m = 4.4 / 1.853 = 2.37 m (solution 1)

for solution 2 5.8 / 1.853 = 3.13 m